# Some explicit values of a $q$-multiple zeta-star function at roots of unity

### v1

@@ 1. Introduction
Then we show some fundamental way to find the values, and give explicit expressions of the values of $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ when $m$ or $s$ are comperatively small. When $n\to\infty$, the values in (1) may be called one of the $q$-multiple zeta-star values (see, e.g., [16, 18]).
@@ 2. $q$-generalized $(r,s)$-Stirling numbers
When $r=1$ and $q\to 1$, $\left\{\!\!\left\{n\atop k\right\}\!\!\right\}^{(s)}=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(1,s)}$ are the Stirling numbers of the second kind with higher level, studied in [12]. When $r=s=1$, $S_{q}(n,k)=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}=(-1)^{n-k}\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(1,1)}$ are the signed $q$-Stirling numbers of the second kind (see, e.g., [5]). When $r=s=1$ and $q\to 1$, $\left\{n\atop k\right\}=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{1}^{(1,1)}$ are the classical Stirling numbers of the second kind.
@@
From Lemma 1 and Lemma 2, $q$-multiple zeta functions and $q$-multiple zeta-star function can be written in terms of $q$-generalized $(1,s)$-Stirling numbers of the first kind and of the second kind, respectively.
@@ 3. The main result
In order to prove Theorem 1, we need the following Lemmas ([6, Proposition 1],[14, Ch.I§2]). Notice that in the case of (2), the equivalent of elementary symmetric functions arises, whereas in the case of (1), the equivalent of complete homogeneous symmetric functions arises.
@@
$\sum_{1\leq j_{1}\leq j_{2}\leq\dots\leq j_{K}\leq M}a_{j_{1}}a_{j_{2}}\cdots a_{j_{K}}=\frac{1}{K!}\mathbf{Y}_{n}(g_{1},1!g_{2},2!g_{3},3!g_{4},\dots)\,,$
@@
  3. (3)
    $a_{n}=\left|\begin{array}[]{ccccc}b_{1}&-1&0&\cdots&0\\ 2b_{2}&b_{1}&-1&&\vdots\\ \vdots&&\ddots&&0\\ (n-1)b_{n-1}&b_{n-2}&\cdots&b_{1}&-1\\ nb_{n}&b_{n-1}&\cdots&b_{2}&b_{1}\\ \end{array}\right|$
  4. (4)
@@
  5. (5)
    $a_{n}=\sum_{j=1}^{n-1}b_{j}a_{n-j}+nb_{n}$
Remark. The relations between (1) and (2) are yielded from a simple modification of Trudi’s formula ([15, Vol.3, p.214],[20]). By applying the inversion formula (see, e.g., [8, Lemma 1],[10, Theorem 1]), we can find that (4) and (5) are equivalent. Others can be entiled by the expansions of the determinants.
@@ 4.3. The cases for $s\geq 3$
Though it becomes more complicated, using $F_{1,1}(X,Y)$, or $F_{2,1}(X,Y)$ and $F_{2,2}(X,Y)$, we can find the expression $\prod_{i=1}^{s}(\alpha_{i}^{n}-1S)$ for $s=1$ or $s=2$, respectively.

### v2

@@ 1. Introduction
Then we show some fundamental ways to find the values, and give explicit expressions of the values of $\mathfrak{Z}_{n}^{\star}(\zeta_{n};m,s)$ when $m$ or $s$ are comparatively small. When $n\to\infty$, the values in (1) may be called one of the $q$-multiple zeta-star values (see, e.g., [17, 19]).
@@ 2. $q$-generalized $(r,s)$-Stirling numbers
When $r=1$ and $q\to 1$, $\left\{\!\!\left\{n\atop k\right\}\!\!\right\}^{(s)}=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{1}^{(1,s)}$ are the Stirling numbers of the second kind with higher level, studied in [13]. When $r=s=1$, $S_{q}(n,k)=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}=(-1)^{n-k}\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{q}^{(1,1)}$ are the signed $q$-Stirling numbers of the second kind (see, e.g., [5]). When $r=s=1$ and $q\to 1$, $\left\{n\atop k\right\}=\left\{\!\!\left\{n\atop k\right\}\!\!\right\}_{1}^{(1,1)}$ are the classical Stirling numbers of the second kind.
@@
From Lemma 1 and Lemma 2, $q$-multiple zeta functions and $q$-multiple zeta-star functions can be written in terms of $q$-generalized $(1,s)$-Stirling numbers of the first kind and of the second kind, respectively.
@@ 3. The main result
In order to prove Theorem 1, we need the following Lemmas ([6, Proposition 1],[15, Ch.I§2]). Notice that in the case of (2), the equivalent of elementary symmetric functions arises, whereas in the case of (1), the equivalent of complete homogeneous symmetric functions arises.
@@
$\sum_{1\leq j_{1}\leq j_{2}\leq\dots\leq j_{K}\leq M}a_{j_{1}}a_{j_{2}}\cdots a_{j_{K}}=\frac{1}{K!}\mathbf{Y}_{K}(g_{1},1!g_{2},2!g_{3},3!g_{4},\dots)\,,$
@@
  3. (iii
    $a_{n}=(-1)^{n-1}\left|\begin{array}[]{ccccc}b_{1}&1&0&\cdots&0\\ 2b_{2}&b_{1}&1&&\vdots\\ \vdots&&\ddots&&0\\ (n-1)b_{n-1}&b_{n-2}&\cdots&b_{1}&1\\ nb_{n}&b_{n-1}&\cdots&b_{2}&b_{1}\\ \end{array}\right|$
  4. (iv)
@@
  5. (v)
    $a_{n}=-\sum_{j=1}^{n-1}b_{j}a_{n-j}+nb_{n}$
Remark. The relations between (i) and (ii) are yielded from a simple modification of Trudi’s formula ([16, Vol.3, p.214],[21]). By applying the inversion formula (see, e.g., [9, Lemma 1],[11, Theorem 1]), we can find that (iv) and (v) are equivalent. Others can be yielded by the expansions of the determinants.
@@ 4.3. The cases for $s\geq 3$
Though it becomes more complicated, using $F_{1,1}(X,Y)$, or $F_{2,1}(X,Y)$ and $F_{2,2}(X,Y)$, we can find the expression $\prod_{i=1}^{s}(\alpha_{i}^{n}-1)$ for $s=1$ or $s=2$, respectively.